OpenStudy (anonymous):

simplify the square root of 25/128

OpenStudy (anonymous):

\[128 = 64 \times 2 = 8^2 \times 2 \]

OpenStudy (anonymous):

yes I know that...I just need to know how to simplify the fraction... 0.o I'm confused and ignorant

OpenStudy (anonymous):

Fine, I'll let you off, this time: \[\sqrt{\frac{25}{128}} = \frac{\sqrt{25}}{\sqrt{64 \times 2}} = \frac{5}{8\sqrt{2}}\] But there isone more step you have to find

OpenStudy (anonymous):

I know what that is and thanks!